OpenStudy (anonymous):

The tangent of the line of the curve y=x^3-6x^2-34x-9

OpenStudy (anonymous):

Did you remember how I showed you?

OpenStudy (anonymous):

sorry I didn't finish the question. It should be: The tangent of the line of the curve y=x^3-6x^2-34x-9 has a slope 2 at two points on the curve. find the two points.

OpenStudy (anonymous):

This one is a tad bit trickier

OpenStudy (anonymous):

y' = 3x^2 - 12x - 34 2 = 3x^2 - 12x - 34 0 = 3x^2 - 12x - 32 0 = 3(x^2 - 4x - 8) 0 = x^2 - 4x - 8 Solve for the zeroes and those are the x values for which there is a slope of 2